A 120-V (line-neutral) wye-connected 4-pole 60-Hz three-phase induction motor is operating at 100% of its 10-kW rated output shaft power at 5% slip. What is the frequency of the rotor electrical currents in Hz?
Answer #1

VL = 120

P = 4  

f = 60

PL = 100 kw

S = 5%

Frequency of rotar fr = S*f

fr = 5/100 * 60

= 0.05*60

fr = 3 Hz